OpenStudy (anonymous):

i need help whats the slope of the line for 2,5 and 4,7

OpenStudy (jdoe0001):

\(\large \bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(2\quad ,&5)\quad &(4\quad ,&7) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (jdoe0001):

ohh ahemm just the slope... .... well then \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(2\quad ,&5)\quad &(4\quad ,&7) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \) :)